OpenStudy (aonz):

Halp please. Medal+fan to best answer If x=log2 and y=log3 express the following in terms of x and y.... log15

OpenStudy (anonymous):

Are you sure this is the correct problem? Because, as stated, the problem doesn't really work.

OpenStudy (aonz):

yea sure is right

OpenStudy (aonz):

i cant figure it out :P

OpenStudy (aonz):

is it x^4 - 1 thats all i can get

OpenStudy (anonymous):

It doesn't work like that. Things inside the log work with multiplication, but since you don't have a factor of 5 anywhere, you can't really make this work.

OpenStudy (aonz):

well in the back of my maths textbook seems very wrong in what it says(to me) y-x+1

OpenStudy (anonymous):

That works lol. I wasn't sure you could use constants.

OpenStudy (aonz):

how does that work?

OpenStudy (anonymous):

\[ \log(15) \equiv \log(\frac{30}{2}) \equiv \log(30) - \log(2) \equiv \log(10 * 3) - x \equiv \log(10) + log(3) - x \equiv 1 + y -x\]

OpenStudy (anonymous):

The cut-off part is the answer 1 + y - x

OpenStudy (aonz):

oh :) thanks so much ill attempted the remainling questions i got stuck on